What is the soultion of the following system of linear equations?
x+2y=5
3x-15=-6y

Answer by tutorcecilia(2152)

x+2y=5
3x-15=-6y
Solve using any method (substitution, elimination, etc.]
.
x+2y=5 [To use substitution, solve for one variable]
y=-1/2x+5/2 [Insert into the other equation and solve for x]
.
3x-15=-6y
3x-15=-6(-1/2x+5/2) [simplify]
3x-15=3x-15
0=0 [The solution is that both equations represent the same line]